OpenStudy (anonymous):

write the equation of the line through the indicated point with the indicated slope. write the final answer in the form y=mx+b m=9/7;(7,-2)

OpenStudy (anonymous):

y=mx+b -2=9/7(7)+b -2=9+b -11=b So the equation is : y=9/7x-11

OpenStudy (anonymous):

would it be x for the whole equation or just the 7

OpenStudy (anonymous):

What do you mean ?

OpenStudy (anonymous):

- it's 9/7 (x)

OpenStudy (anonymous):

ok